# A Survey of Slow Thinking-based Reasoning LLMs using Reinforced Learning and Inference-time Scaling Law

### v1

@@ Abstract.
This survey explores recent advancements in reasoning large language models (LLMs) designed to mimic “slow thinking”—a reasoning process inspired by human cognition, as described in Kahneman’s Thinking, Fast and Slow. These models, like OpenAI’s o1, focus on scaling computational resources dynamically during complex tasks, such as math reasoning, code generation, and agents. We present the development of reasoning LLMs and list their key technologies. By synthesizing over 100 studies, it charts a path toward LLMs that combine human-like deep thinking with scalable efficiency for reasoning. The review breaks down methods into three categories: (1) test-time scaling dynamically adjusts computation based on task complexity via search and sampling, dynamic verification; (2) reinforced learning refines decision-making through iterative improvement leveraging policy networks, reward models, and self-evolution strategies; and (3) slow-thinking frameworks (e.g., long CoT, hierarchical processes, hybrid thinking) that structure problem-solving with manageable steps. The survey highlights the challenges and further directions of this domain. Understanding and advancing the reasoning abilities of LLMs is crucial for unlocking their full potential in real-world applications, from scientific discovery to decision support systems.
@@ 1. Introduction
This survey aims to synthesize insights from over 100 studies to chart the evolution of slow-thinking-based reasoning LLMs. It categorizes current methodologies into three key areas: (1) test-time scaling, which adjusts computational effort based on task complexity; (2) reinforced learning, which enhances decision-making through reward models and self-improvement strategies; and (3) slow-thinking frameworks, which structure problem-solving using techniques such as long chain-of-thought (CoT) and hierarchical reasoning processes. Through this exploration, we aim to provide a comprehensive roadmap for future research in slow-thinking-based LLMs.
@@ 2. The Development of Reasoning LLMs
Recently, o1-like models have emerged as a family of reasoning language models (RLMs) designed to replicate or extend the advanced reasoning, search, and alignment capabilities pioneered by OpenAI’s o1 series (Zeng et al., 2024a; OpenAI, 2024) (Fig. LABEL:fig:_timeline). These models typically share several core design principles. First, they incorporate reinforcement learning (RL) to optimize policy behaviors for complex reasoning tasks (El-Kishky et al., 2025; Guo et al., 2025; Team et al., 2025). This often takes the form of Process Reward Models (PRMs) and Outcome Reward Models (ORMs), which respectively evaluate intermediate reasoning steps and final answers (Lightman et al., 2023; Guo et al., 2025; Xu et al., 2024a). Second, they emphasize long COT or slow thinking paradigms, allowing the model to reason in multiple stages, verify partial solutions, and refine its outputs via techniques such as self-verification or guided search (Guo et al., 2025; Kumar et al., 2025; Xu et al., 2024a). Third, search-based methods—for instance, beam search, Monte Carlo Tree Search (MCTS), or retrieval-augmented generation—serve as additional mechanisms to explore and validate candidate reasoning paths before finalizing an answer (El-Kishky et al., 2025; Zhao et al., [n. d.]; Zhang et al., 2024d; Zhong et al., 2024).
Beyond these core techniques, many o1-like frameworks introduce multi-stage training pipelines that blend supervised fine-tuning (SFT) on human-curated CoT data with RL-driven policies for iterative refinement (El-Kishky et al., 2025; Guo et al., 2025; Team et al., 2025; Kumar et al., 2025). This often includes cold-start or rule-based reward functions to bootstrap reasoning patterns, which are then iteratively replaced by more advanced reward models once the base model attains a certain level of proficiency (Guo et al., 2025; Team et al., 2025). Active learning strategies, where high-uncertainty or error-prone samples are re-labeled or refined, further enhance data efficiency and model alignment (Lightman et al., 2023; Xu et al., 2024a). Some approaches also incorporate specialized modules—for example, retrieval-augmented generation (RAG) or verifiers—to manage external retrieval and mitigate hallucinations (Li et al., 2025a; Xu et al., 2024a; Zhao et al., [n. d.]; Zhang et al., 2024d).
@@ 3.6. Self-training
Search [ E.g., Marco-o1 (Zhao et al., [n. d.]), STILL-1 (Jiang et al., 2024), REBASE (Wu et al., 2024c), ToT (Yao et al., 2023) , leaf, text width=44.6em ] ] [ Sampling [ E.g., Brown et al. (Brown et al., 2024), SANA1.5 (Xie et al., 2025), Speculative Rejection (Sun et al., 2024a), S1 (Muennighoff et al., 2025), Feynman-Kac Steering (Singhal et al., 2025) , leaf, text width=44.6em ] ] [ Long to Short CoT [ E.g., OverThink (Kumar et al., 2025), LightThinker (Zhang et al., 2025d), TokenSkip (Xia et al., 2025), TALE (Han et al., 2024),
COD (Xu et al., 2025b), InftyThink (Yan et al., 2025), SoT (Aytes et al., 2025) , leaf, text width=44.6em ] ] ] [ Dynamic Verification
@@
Backtracking. [ E.g., Deepseek-R1 (Guo et al., 2025), RedStar (Xu et al., 2025a), Distilling System 2 into System 1 (Yu et al., 2024),
@@
6.2) [ Explicit Structures [ E.g., ReasonFlux (Yang et al., 2025), Li et al. (Li et al., 2025a) ,leaf, text width=44.6em ] ] [ Agentic Systems [ E.g., MALT (Motwani et al., 2025), OctoTools (Lu et al., 2025a), Agentic Reasoning (Wu et al., 2025) ,leaf, text width=44.6em ] ] [ Dynamic Control Mechanism [ E.g., MixLLM (Wang et al., 2025), AdaptiveStep (Liu et al., 2025b) ,leaf, text width=44.6em ] ] [ Latent Space Manipulations [ E.g., Yang et al. (Yang et al., 2023), LTMs (Kong et al., 2025), PEARL (Chen et al., 2025b), ThoughtProbe (Wang and Xu, 2025) ,leaf, text width=44.6em ] ] ] [ Hybrid Thinking
@@
Architectures [ E.g., Talker-Reasoner (Christakopoulou et al., 2024), FS-GEN (Zhang et al., 2024b), SYMBOLIC-MoE (Chen et al., 2025c), Lemmanaid (Alhessi et al., 2025) ,leaf, text width=44.6em ] ] [ Tailored Training [ E.g., RELAY (Yu et al., 2025), Mix Distillation (Li et al., 2025b), MoBA (Lu et al., 2025b), B-STaR (Zeng et al., 2025c), FAST (Sun et al., 2024b), Hu et al. (Hu et al., 2023),
@@ 4.1.3. Long to Short CoT
Efforts to optimize long COT reasoning in LLMs have led to diverse innovative approaches aimed at improving efficiency while maintaining reasoning performance. The OverThink framework (Kumar et al., 2025) highlights vulnerabilities in reasoning LLMs, demonstrating how slowdown attacks can disrupt inference efficiency by injecting decoy reasoning problems. To counter such inefficiencies, methods like LightThinker propose dynamically compressing intermediate reasoning steps, achieving faster inference on complex tasks with minimal performance trade-offs (Zhang et al., 2025d). Similarly, the TokenSkip strategy (Xia et al., 2025) enables selective skipping of less critical tokens, providing controllable CoT compression, while a token-budget-aware framework adjusts the number of reasoning tokens based on task complexity, optimizing resource allocation (Han et al., 2024). Other frameworks, such as Chain of Draft (Xu et al., 2025b), focus on generating concise yet informative intermediate outputs to accelerate reasoning, and the InftyThink paradigm (Yan et al., 2025) restructures reasoning into iterative processes with intermediate summaries, significantly reducing computational demands. Additionally, the Sketch-of-Thought (SoT) framework (Aytes et al., 2025) integrates cognitive-inspired strategies with linguistic constraints to minimize token usage while preserving accuracy. Collectively, these approaches address the challenges of long CoT reasoning, enabling more efficient and scalable reasoning systems for practical applications.
@@ 6.1.4. Reflection and Backtracking
Reflection and backtracking mechanisms facilitate a model’s ability to monitor internal reasoning processes, detect errors, and dynamically adjust its reasoning trajectories. Numerous studies have incorporated techniques for overseeing intermediate reasoning steps. For instance, Guo et al. (Guo et al., 2025) proposed a Self-Refinement mode wherein the model continuously evaluates and, if necessary, corrects intermediate outputs via recursive checks. Similarly, Xu et al. (Xu et al., 2025a) presented the RedStar framework, which employs a ”step-by-step verification” process to identify and flag errors within the generated reasoning chain, although its reported efficacy can be context-dependent. Furthermore, methods developed by Yu et al. (Yu et al., 2024) and Qi et al. (Qi et al., 2024) specifically emphasize backtracking capabilities and associated techniques like gradient stabilization. Upon detecting a discrepancy during a reasoning step, the model can re-evaluate prior assumptions and adjust subsequent steps accordingly. For example, in a complex mathematical derivation, identification of an error at a specific computational step allows the model to backtrack, recalculate, and revise subsequent steps. Moreover, Min et al. (Min et al., 2024b) described a self-improvement paradigm where the model iteratively generates high-quality reasoning demonstrations, which are then incorporated into its training data. This continuous learning mechanism enables the model to progressively refine its reasoning strategies. Finally, Gan et al. (Gan et al., 2025) employed tree search algorithms, such as MCTS, to explore multiple potential reasoning paths. Reward models were subsequently utilized to select the optimal path, thereby effectively reflection and backtracking into the model’s decision-making framework.
@@ 6.2. Hierarchical Reasoning
Innovations increasingly target the model’s internal processes and representations. (Hao et al., 2024) Strategies include iterative refinement for enhanced In-Context Learning (Yang et al. (Yang et al., 2023), the introduction of explicit latent thought vectors for modular control (Kong et al., 2025), adversarial training frameworks for intrinsic permutation robustness (Chen et al., 2025b), and classifier-guided exploration of latent reasoning pathways (Wang and Xu, 2025).
@@ 6.3. Hybrid Thinking
HTM principles also manifest in specialized architectures and collaborative model setups, demonstrating structural ways to embody the dual-process approach. Examples include agent systems with distinct “Talker” (fast) and “Reasoner” (slow) roles (Christakopoulou et al., 2024), collaborations between large (slow) and small (fast) models like FS-GEN (Zhang et al., 2024b), skill-based Mixture-of-Experts routing (SYMBOLIC-MoE (Chen et al., 2025c)), and neuro-symbolic tools like Lemmanaid (Alhessi et al., 2025) combining fast neural generation with slow symbolic validation.

### v2

@@ Abstract.
This survey explores recent advancements in reasoning large language models (LLMs) designed to mimic “slow thinking”—a reasoning process inspired by human cognition, as described in Kahneman’s Thinking, Fast and Slow. These models, like OpenAI’s o1, focus on scaling computational resources dynamically during complex tasks, such as math reasoning, code generation, and agents. We present the development of reasoning LLMs and list their key technologies. By synthesizing over 160 studies, it charts a path toward LLMs that combine human-like deep thinking with scalable efficiency for reasoning. The review breaks down methods into three categories: (1) test-time scaling dynamically adjusts computation based on task complexity via search and sampling, dynamic verification; (2) reinforced learning refines decision-making through iterative improvement leveraging policy networks, reward models, and self-evolution strategies; and (3) slow-thinking frameworks (e.g., long CoT, hierarchical processes, hybrid thinking) that structure problem-solving with manageable steps. The survey highlights the challenges and further directions of this domain. Understanding and advancing the reasoning abilities of LLMs is crucial for unlocking their full potential in real-world applications, from scientific discovery to decision support systems.
@@ 1. Introduction
Recent advancements in the complex reasoning capabilities of LLMs have catalyzed substantial survey efforts within the academic community. These studies analyze the burgeoning field of reasoning LLMs from various perspectives: some concentrate on particular dimensions, such as the features and mechanisms of Long Chain-of-Thought (Chen et al., 2025b; Liu et al., 2023); others emphasize core technical approaches, particularly positioning Reinforcement Learning (RL) as a primary driver for advancing model capabilities (Zeng et al., 2024a; Xu et al., 2025a); certain works analyze model advancement via specific training paradigms, such as self-evolution (He et al., 2025; Tao et al., 2024); while others offer comprehensive overviews of technical stages like post-training, including fine-tuning, RL, and test-time scaling (Kumar et al., 2025a). Notably, reference (Li et al., 2025c) aligns conceptually with the present work by also leveraging the dual-system theory (System 1/System 2) from cognitive science as an analytical framework.
Positioned within this context, the present survey provides a unique and focused perspective. Our primary contribution centers on the analysis of the cognitive goal of “Slow Thinking” (analogous to System 2), distinctively identifying RL alongside the Inference-time Scaling Law as the principal, synergistic technical mechanisms for achieving this advanced reasoning capability. Although RL and test-time scaling are addressed individually or within wider scopes in other surveys (Zeng et al., 2024a; Xu et al., 2025a; Li et al., 2025c), this review uniquely synthesizes these two methodologies within the “Slow Thinking” conceptual framework through systematic integration and in-depth analysis. Particularly, this survey aims to synthesize insights from over 160 studies to chart the evolution of slow-thinking-based reasoning LLMs. It categorizes current methodologies into three key areas: (1) test-time scaling, which adjusts computational effort based on task complexity; (2) reinforced learning, which enhances decision-making through reward models and self-improvement strategies; and (3) slow-thinking frameworks, which structure problem-solving using techniques such as long chain-of-thought (CoT) and hierarchical reasoning processes. Through this exploration, we aim to provide a comprehensive roadmap for future research in slow-thinking-based LLMs.
@@ 2. The Development of Reasoning LLMs
Recently, o1-like models have emerged as a family of reasoning language models (RLMs) designed to replicate or extend the advanced reasoning, search, and alignment capabilities pioneered by OpenAI’s o1 series (Zeng et al., 2024a; OpenAI, 2024) (Fig. LABEL:fig:_timeline). These models typically share several core design principles. First, they incorporate reinforcement learning (RL) to optimize policy behaviors for complex reasoning tasks (El-Kishky et al., 2025; Guo et al., 2025; Team et al., 2025). This often takes the form of Process Reward Models (PRMs) and Outcome Reward Models (ORMs), which respectively evaluate intermediate reasoning steps and final answers (Lightman et al., 2023; Guo et al., 2025; Xu et al., 2024a). Second, they emphasize long COT or slow thinking paradigms, allowing the model to reason in multiple stages, verify partial solutions, and refine its outputs via techniques such as self-verification or guided search (Guo et al., 2025; Kumar et al., 2025b; Xu et al., 2024a). Third, search-based methods—for instance, beam search, Monte Carlo Tree Search (MCTS), or retrieval-augmented generation—serve as additional mechanisms to explore and validate candidate reasoning paths before finalizing an answer (El-Kishky et al., 2025; Zhao et al., [n. d.]; Zhang et al., 2024d; Zhong et al., 2024).
Beyond these core techniques, many o1-like frameworks introduce multi-stage training pipelines that blend supervised fine-tuning (SFT) on human-curated CoT data with RL-driven policies for iterative refinement (El-Kishky et al., 2025; Guo et al., 2025; Team et al., 2025; Kumar et al., 2025b). This often includes cold-start or rule-based reward functions to bootstrap reasoning patterns, which are then iteratively replaced by more advanced reward models once the base model attains a certain level of proficiency (Guo et al., 2025; Team et al., 2025). Active learning strategies, where high-uncertainty or error-prone samples are re-labeled or refined, further enhance data efficiency and model alignment (Lightman et al., 2023; Xu et al., 2024a). Some approaches also incorporate specialized modules—for example, retrieval-augmented generation (RAG) or verifiers—to manage external retrieval and mitigate hallucinations (Li et al., 2025a; Xu et al., 2024a; Zhao et al., [n. d.]; Zhang et al., 2024d).
@@ 3.6. Self-training
Search [ E.g., Marco-o1 (Zhao et al., [n. d.]), STILL-1 (Jiang et al., 2024), REBASE (Wu et al., 2024c), ToT (Yao et al., 2023) , leaf, text width=44.6em ] ] [ Sampling [ E.g., Brown et al. (Brown et al., 2024), SANA1.5 (Xie et al., 2025), Speculative Rejection (Sun et al., 2024a), S1 (Muennighoff et al., 2025), Feynman-Kac Steering (Singhal et al., 2025) , leaf, text width=44.6em ] ] [ Long to Short CoT [ E.g., OverThink (Kumar et al., 2025b), LightThinker (Zhang et al., 2025d), TokenSkip (Xia et al., 2025), TALE (Han et al., 2024),
COD (Xu et al., 2025c), InftyThink (Yan et al., 2025), SoT (Aytes et al., 2025) , leaf, text width=44.6em ] ] ] [ Dynamic Verification
@@
Backtracking. [ E.g., Deepseek-R1 (Guo et al., 2025), RedStar (Xu et al., 2025b), Distilling System 2 into System 1 (Yu et al., 2024),
@@
6.2) [ Explicit Structures [ E.g., ReasonFlux (Yang et al., 2025), Li et al. (Li et al., 2025a) ,leaf, text width=44.6em ] ] [ Agentic Systems [ E.g., MALT (Motwani et al., 2025), OctoTools (Lu et al., 2025a), Agentic Reasoning (Wu et al., 2025) ,leaf, text width=44.6em ] ] [ Dynamic Control Mechanism [ E.g., MixLLM (Wang et al., 2025), AdaptiveStep (Liu et al., 2025b) ,leaf, text width=44.6em ] ] [ Latent Space Manipulations [ E.g., Yang et al. (Yang et al., 2023), LTMs (Kong et al., 2025), PEARL (Chen et al., 2025c), ThoughtProbe (Wang and Xu, 2025) ,leaf, text width=44.6em ] ] ] [ Hybrid Thinking
@@
Architectures [ E.g., Talker-Reasoner (Christakopoulou et al., 2024), FS-GEN (Zhang et al., 2024b), SYMBOLIC-MoE (Chen et al., 2025d), Lemmanaid (Alhessi et al., 2025) ,leaf, text width=44.6em ] ] [ Tailored Training [ E.g., RELAY (Yu et al., 2025), Mix Distillation (Li et al., 2025b), MoBA (Lu et al., 2025b), B-STaR (Zeng et al., 2025c), FAST (Sun et al., 2024b), Hu et al. (Hu et al., 2023),
@@ 4.1.3. Long to Short CoT
Efforts to optimize long COT reasoning in LLMs have led to diverse innovative approaches aimed at improving efficiency while maintaining reasoning performance. The OverThink framework (Kumar et al., 2025b) highlights vulnerabilities in reasoning LLMs, demonstrating how slowdown attacks can disrupt inference efficiency by injecting decoy reasoning problems. To counter such inefficiencies, methods like LightThinker propose dynamically compressing intermediate reasoning steps, achieving faster inference on complex tasks with minimal performance trade-offs (Zhang et al., 2025d). Similarly, the TokenSkip strategy (Xia et al., 2025) enables selective skipping of less critical tokens, providing controllable CoT compression, while a token-budget-aware framework adjusts the number of reasoning tokens based on task complexity, optimizing resource allocation (Han et al., 2024). Other frameworks, such as Chain of Draft (Xu et al., 2025c), focus on generating concise yet informative intermediate outputs to accelerate reasoning, and the InftyThink paradigm (Yan et al., 2025) restructures reasoning into iterative processes with intermediate summaries, significantly reducing computational demands. Additionally, the Sketch-of-Thought (SoT) framework (Aytes et al., 2025) integrates cognitive-inspired strategies with linguistic constraints to minimize token usage while preserving accuracy. Collectively, these approaches address the challenges of long CoT reasoning, enabling more efficient and scalable reasoning systems for practical applications.
@@ 6.1.4. Reflection and Backtracking
Reflection and backtracking mechanisms facilitate a model’s ability to monitor internal reasoning processes, detect errors, and dynamically adjust its reasoning trajectories. Numerous studies have incorporated techniques for overseeing intermediate reasoning steps. For instance, Guo et al. (Guo et al., 2025) proposed a Self-Refinement mode wherein the model continuously evaluates and, if necessary, corrects intermediate outputs via recursive checks. Similarly, Xu et al. (Xu et al., 2025b) presented the RedStar framework, which employs a ”step-by-step verification” process to identify and flag errors within the generated reasoning chain, although its reported efficacy can be context-dependent. Furthermore, methods developed by Yu et al. (Yu et al., 2024) and Qi et al. (Qi et al., 2024) specifically emphasize backtracking capabilities and associated techniques like gradient stabilization. Upon detecting a discrepancy during a reasoning step, the model can re-evaluate prior assumptions and adjust subsequent steps accordingly. For example, in a complex mathematical derivation, identification of an error at a specific computational step allows the model to backtrack, recalculate, and revise subsequent steps. Moreover, Min et al. (Min et al., 2024b) described a self-improvement paradigm where the model iteratively generates high-quality reasoning demonstrations, which are then incorporated into its training data. This continuous learning mechanism enables the model to progressively refine its reasoning strategies. Finally, Gan et al. (Gan et al., 2025) employed tree search algorithms, such as MCTS, to explore multiple potential reasoning paths. Reward models were subsequently utilized to select the optimal path, thereby effectively reflection and backtracking into the model’s decision-making framework.
@@ 6.2. Hierarchical Reasoning
Innovations increasingly target the model’s internal processes and representations. (Hao et al., 2024) Strategies include iterative refinement for enhanced In-Context Learning (Yang et al. (Yang et al., 2023), the introduction of explicit latent thought vectors for modular control (Kong et al., 2025), adversarial training frameworks for intrinsic permutation robustness (Chen et al., 2025c), and classifier-guided exploration of latent reasoning pathways (Wang and Xu, 2025).
@@ 6.3. Hybrid Thinking
HTM principles also manifest in specialized architectures and collaborative model setups, demonstrating structural ways to embody the dual-process approach. Examples include agent systems with distinct “Talker” (fast) and “Reasoner” (slow) roles (Christakopoulou et al., 2024), collaborations between large (slow) and small (fast) models like FS-GEN (Zhang et al., 2024b), skill-based Mixture-of-Experts routing (SYMBOLIC-MoE (Chen et al., 2025d)), and neuro-symbolic tools like Lemmanaid (Alhessi et al., 2025) combining fast neural generation with slow symbolic validation.